OpenStudy (anonymous):

Help ! i am getting really confused reading this!!!! Tabitha is renting a ballroom for an upcoming dinner reception. The rental fee for the ballroom includes a base price plus the cost of dinner per person. The table below shows the cost of the ballroom rental in relation to the number of people attending the dinner reception. No. of Attendees 15 30 45 Total Rental Cost $375 $480 $585 What is the base price for the ballroom rental? A. $270 B. $7 C. $105 D. $54

OpenStudy (anonymous):

I'm confused too omfg

OpenStudy (anonymous):

cost = base + (number of people)(price per person) use their numbers to find base. get cost per person from change in cost/change in # people

OpenStudy (anonymous):

true @shoplovepc lol

OpenStudy (anonymous):

I think I get it!

OpenStudy (anonymous):

and what do you mean?@douglaswinslowcooper

OpenStudy (anonymous):

the price goes UP by 105 per 15 extra people

OpenStudy (anonymous):

oh wow your right lol

OpenStudy (anonymous):

@douglaswinslowcooper is that right?

OpenStudy (anonymous):

wait its asking what the base price for the ball room

OpenStudy (anonymous):

Thus cost = base + (number of people)($7/person) so 375 = base + 15(7) = base + 105 base = 375-105 = 270 A is answer.

OpenStudy (anonymous):

@douglaswinslowcooper that makes more sense @Louis_Tomlinson_Girl nvm I was wrong lol

OpenStudy (anonymous):

wow that was a long question thanks to both of you ( i wish i could give out 2 metals!